OpenStudy (zzr0ck3r):

Let G be a group Let H and K be subgroups of G with H normal in G define HK:= {xy|x in G ^ y in K} prove H is a normal subgroup of HK

OpenStudy (zzr0ck3r):

proof: Let h be in H then (xy)*h(xy)^-1 = xyhy^(-1)x^(-1) = xh'x^(-1) is in H this is because yhy^(-1) is in H because y is in G and xhx^(-1) is in H because x is in G

OpenStudy (zzr0ck3r):

this look ok?

OpenStudy (nincompoop):

the format doesn't fit the conventional notation of providing a proof :( there's a proof wiki, see if that'll help :)

OpenStudy (nincompoop):

found it http://www.proofwiki.org/wiki/Definition:Normal_Subgroup

OpenStudy (zzr0ck3r):

look at definition 3

OpenStudy (zzr0ck3r):

or better yet look at the also known as...

OpenStudy (kinggeorge):

Your proof looks fine to me. The one thing I have to add, is that you may need to show that \(HK\) is a subgroup of \(G\). In fact, this is true if \(HK\) is a subset of the normalizer of \(H\) in \(G\). I.e., if \(H\lhd G\), then \(HK\) is a subgroup of \(G\).